OpenStudy (anonymous):

select the pair of functions representing f(x) and g(x) if f(g(x))=18x-1. Choices are f(x)=3x+2 or 3x-2 or 3x-5 or 3x+5

OpenStudy (ranga):

which ones are the g(x)?

OpenStudy (anonymous):



OpenStudy (anonymous):

I don't know. that is all of the problem. It is missing something isn't it?

OpenStudy (ranga):

Yes. You will need f(x) and g(x). See if this question applies to a previous problem or makes reference something else on that page or chapter.

OpenStudy (anonymous):

It doesn't go with any other problem. I thought something was missing. I will chck with the teacher . thanks

OpenStudy (ranga):

See if you can take a picture of it and post it here. I will see if I can spot the problem.

OpenStudy (anonymous):

I typed it exactly as given to me.

OpenStudy (ranga):

I am more interested in the answer choices because f(x) and g(x) are both in each choice of answers. I don't think you typed it exactly as it appears because I don't think it has: f(x)=3x+2 or 3x-2 or 3x-5 or 3x+5 all in one line like that.

OpenStudy (anonymous):

f(x)=3x+2 f(x)=3x-5 f(x)=3x-2 f(x)=3x+5

OpenStudy (ranga):

There must be one more list that has 4 choices for g(x) g(x) = g(x) = g(x) = g(x) = The g(x) may even be defined earlier for a previous problem.

OpenStudy (anonymous):

nope

OpenStudy (ranga):

or at least one g(x) somewhere.

OpenStudy (anonymous):

i will check with teacher

OpenStudy (ranga):

okay.